Please can you help me solve x
(((3 + 3/x = 7)))
Thank you

Answer by AnlytcPhil(1806)

Please can you help me solve x 


            3 + 3/x = 7 

The LCD = x, so multiply every term by 

        3 + 3/x = 7

              1 
        3 + x·3/x = 7
                  1

             3x + 3 = 7x

Add -3 to both sides

                 3x = 7x - 3

Add -7x to both sides

                -4x = -3

Divide both sides by -4

                  x = -3/(-4)

                  x = 3/4

Edwin